A 2L closed vessel contain 0.1mol nitrogen and 0.3 mol hydrogen, the gas reacted to form ammonia gas. 10min later, the vessel contain 0.22mol hydrogen gas. What is the average rate of hydrogen consume in molL-1min-1.
1
Expert's answer
2020-04-06T11:46:43-0400

The average rate of gas consumption is :

"r = \frac{n}{V\cdot t}" ,

where "n" is the number of the moles of the gas, "V" is the volume of the vessel and "t" is the time elapsed.

In this case, the rate of hydrogen consumption is:

"r = \frac{0.30 - 0.22}{2\cdot10} = 0.0040" mol/L/min, or 4·10-3 mol L-1 min-1

Answer: The average rate of hydrogen consumption is 4·10-3 mol L-1 min-1.